# Semi-Supervised Class-Agnostic Motion Prediction with Pseudo Label Regeneration and BEVMix

### 13 Dec 2023 v1

@@ Introduction
Furthermore, inspired by the local rigid characteristics of most objects (e.g., cars) in the driving scenario, we re-generate pseudo labels for unreliable cells from reliable adjacent labels, as shown in Algorithm 1. For each unreliable cell, we first find $K$ nearest reliable neighbor cells based on Euclidean distance (lines 3-5). And the neighbors are valid only if the distances are within the distance threshold $\beta$ (lines 6-7). Then we use these reliable labels to re-generate labels for unreliable cells. We first calculate the weights $W$ based on the distance, the closer the distance is, the larger the weight is (line 12). Then we calculate the weighted mean of the reliable neighbors as the re-generated labels (line 13). To evaluate the quality of the re-generated labels, We also calculate the weighted differences between the K neighbors and $M_{mean}$ to measure the local consistency $H$ (lines 14-15). The idea is derived from the fact that most objects in the scene are almost rigid and each part of a single one should have similar motion. Therefore, a larger $H$ indicates that the unreliable cell and its K neighbors are more likely from the same object. By discarding re-generated labels with low consistency $H$, we obtain the final pseudo labels $M^{u}$ and correspondent index $I^{u}$ for the unlabeled sample (line 17). During training, we only compute unsupervised loss for cells within index $I^{u}$.
@@
As discussed by FixMatch (Sohn et al. 2020a), weak-to-strong consistency is important for the performance of semi-supervised learning. In our framework, we establish consistency regularization through a two-step process: firstly, the teacher model processes weakly augmented unlabeled samples, generating pseudo-labels through test-time inference. Subsequently, the same samples undergo strong augmentation and are fed to the student model for predictions. Finally, the regularization is achieved by enforcing consistency between the predictions from strongly augmented samples and the pseudo-labels derived from weakly augmented samples. Although data augmentation strategies have been widely studied in image classification, object-detection, etc. (Zhang et al. 2018; Hendrycks et al. 2019; Yun et al. 2019), not yet in class-agnostic motion prediction tasks. To this end, we introduce two simple but effective data augmentations called temporal-sampling and BEVMix.
@@
Temporal-sampling. Taking input $\mathcal{V}=\{V^{t}\}_{t=1}^{T}$, we can generate artificial sequence $\mathcal{V_{TS}}=\{V^{1},V^{3},V^{5},...,V^{t}\}$ by uniform-sampling along temporal dimension. And correspondent motion ground truth will be $M_{TS}=2M$. We then pad $\mathcal{V_{TS}}=\{V^{1},...,V^{1},V^{3},V^{5},...,V^{t}\}$ with $V^{1}$ to make sure the sequence length is equal to $\mathcal{V}$. From temporal sampling, we simulate a scenario from original input $\mathcal{V}$, in which objects are static at the beginning and suddenly start to move with double speed. By using temporal sampling, we can compensate for a number of fast-speed samples, which are relatively fewer in comparison to the static and slow ones in the original dataset.
@@
BEVMix. Mix data augmentations aim to generate artifact samples by combining two existing samples. Mixup (Zhang et al. 2018) proposes to enforce smooth label transitions by linearly interpolating between pairs of input samples and their corresponding class labels. Instead of generating unnatural mixed samples like Mixup, CutMix (Yun et al. 2019) proposes directly replacing the image region with a patch from another training image. These strategies are also lifted from 2D to 3D for point cloud recognition (Chen et al. 2020; Zhang et al. 2022). To be not limited to single objects, Mix3D (Nekrasov et al. 2021) proposes to mix two full point cloud scenes directly by combining all the points from two scenes as a single scene for point cloud segmentation. Although effective, previous data mix approaches mainly focus on single input while the motion prediction methods take sequence input, as it is important to observe the locations of each object throughout the temporal window to make accurate motion prediction. Directly applying data mix strategies like Cutmix may lead to a case in which we cannot find correspondent cells of a moving object in the current frame from previous frames. Additionally, these data mix approaches are mainly designed for classification or segmentation based on some task-specific priors, which are also not suitable for the motion prediction task.

### 14 Dec 2023 v2

@@ Introduction
Furthermore, inspired by local rigid characteristics of most objects (e.g., cars) in the driving scenario, we re-generate pseudo labels for unreliable cells from reliable adjacent labels, as shown in Algorithm 1. For each unreliable cell, we first find $K$ nearest reliable neighbor cells based on Euclidean distance (lines 3-5). And the neighbors are valid only if the distances are within the distance threshold $\beta$ (lines 6-7). Then we use these reliable labels to re-generate labels for unreliable cells. We first calculate the weights $W$ based on the distance, the closer the distance is, the larger the weight is (line 12). Then we calculate the weighted mean of the reliable neighbors as the re-generated labels (line 13). To evaluate the quality of the re-generated labels, We calculate the weighted differences between the K neighbors and $M_{mean}$ to measure the local consistency $H$ (lines 14-15). The idea is derived from the fact that most objects in the scene are almost rigid and each part of a single one should have similar motion. A larger $H$ indicates that the unreliable cell and its K neighbors are more likely from the same object. By discarding re-generated labels with low consistency $H$, we obtain the final pseudo labels $M^{u}$ and correspondent index $I^{u}$ for the unlabeled sample (line 17). We only compute unsupervised loss for cells within index $I^{u}$.
@@
As discussed by FixMatch (Sohn et al. 2020a), weak-to-strong consistency is important for the performance of semi-supervised learning. In our framework, we establish consistency regularization through following process: firstly, the teacher model processes weakly augmented unlabeled samples, generating pseudo-labels through test-time inference. Subsequently, the same samples undergo strong augmentation and are fed to the student model for predictions. Finally, the regularization is achieved by enforcing consistency between the predictions from strongly augmented samples and the pseudo-labels derived from weakly augmented samples. While data augmentation has been widely studied in image classification and object detection, it have not yet been thoroughly explored in the class-agnostic motion prediction task. To this end, we introduce two simple yet effective data augmentations called temporal-sampling and BEVMix.
@@
Temporal-sampling. Taking input $\mathcal{V}=\{V^{t}\}_{t=1}^{T}$, we can generate artificial sequence $\mathcal{V_{TS}}=\{V^{1},V^{3},V^{5},...,V^{t}\}$ by uniform-sampling along temporal dimension. And correspondent motion ground truth will be $M_{TS}=2M$. We then pad $\mathcal{V_{TS}}=\{V^{1},...,V^{1},V^{3},V^{5},...,V^{t}\}$ with $V^{1}$ to make sure the sequence length is equal to $\mathcal{V}$. From temporal sampling, we simulate a scenario from original input $\mathcal{V}$, in which objects are static at the beginning and suddenly start to move with double speed. By temporal-sampling, we can compensate for a number of fast-speed samples, which are relatively fewer in comparison to the static and slow ones in the original dataset.
@@
BEVMix. Mix data augmentations aim to generate artifact samples by combining two existing samples. Mixup (Zhang et al. 2018) proposes to enforce smooth label transitions by linearly interpolating between pairs of input samples and their corresponding class labels. Instead of generating unnatural mixed samples like Mixup, CutMix (Yun et al. 2019) proposes directly replacing the image region with a patch from another training image. These strategies are also lifted from 2D to 3D for point cloud recognition (Chen et al. 2020; Zhang et al. 2022). To be not limited to single objects, Mix3D (Nekrasov et al. 2021) proposes to mix two full point cloud scenes directly by combining all the points from two scenes as a single scene for point cloud segmentation. LaserMix (Kong et al. 2023) leverages the spatial prior of point clouds and mix two scenes by LiDAR scans. Although effective, previous data mix approaches mainly focus on single input while the motion prediction methods take sequence input, as it is important to observe the locations of each object throughout the temporal window to make accurate motion prediction. Directly applying data mix strategies like Cutmix may lead to a case in which we cannot find correspondent cells of a moving object in the current frame from previous frames. Additionally, these data mix approaches are mainly designed for classification or segmentation based on some task-specific priors, which are also not suitable for the motion prediction task.